# Exploring Self-Supervised Vision Transformers for Deepfake Detection: A Comparative Analysis

### v1

@@ 5.2 Approach 2: Fine-tuning final transformer blocks
We selected DINOv2 - ViT-L/14-Reg, extensively studied in Approach 1, as the representative of DINOv2. Similarly, we chose EfficientNetV2, DeiT III, and EVA-CLIP for comparison. The performances of fine-tuning the final block (and also the tokens in the case of transformers) are shown in Table 7. Compared to Approach 1, all models gained better results, and the performance gaps between DINOv2 and others were reduced, with EVA-CLIP being the closest competitor. Nevertheless, DINOv2 remained the top performer. To narrow the gap with DINOv2, EVA-CLIP would need to be pre-trained with a vast dataset featuring rich annotations—a costly endeavor compared to DINOv2, which was pre-trained on a substantially smaller dataset without any annotations. Given the same architecture (DeiT III and DINOv2), the performance gap is nearly 6% in terms of EER. In addition to different training receipts, part of this drop may stem from the larger training data (although without annotations). Overall, these results again underscore the significant advantage of using SSL for pre-training ViTs.
Table 7: Comparison between different ConvNet and transformer architectures on Approach 2. The final block of each pre-trained backbone was fine-tuned for deepfake detection. Accuracies were calculated using a threshold of 0.5.
Architecture Accuracy EER HTER
EfficientNetV2 Large 84.76 15.05 15.21
DeiT III L/16-LayerScale 82.64 17.21 17.28
EVA-02-CLIP-L/14 88.29 11.77 11.77
DINOv2 ViT-L/14-Reg 88.78 11.32 11.26
We next conducted an ablation study to determine the optimal number of $k$ final blocks required for fine-tuning. It is important to note that different DINO backbone sizes have varying numbers of blocks; for example, DINOv2 - ViT-L/14-Reg has 24 blocks. The results are visualized in Fig. 2. If $k$ is small, the model may not have enough parameters to adequately adapt to the new task, resulting in underfitting. Conversely, the model may experience overfitting, particularly when the fine-tuning dataset is small. The optimal number of $k$ is approximately half of the total blocks, which in this experiment is 11. When increasing $k$ from 1 to 11, the EERs decrease from 11.32% to 5.63%, representing an improvement of 5.69%.
Figure 2: Results of ablation study for Approach 2. The numbers of fine-tuned final transformer blocks ($k$) and their corresponding EERs are shown. SSL pre-trained DINOv2 - ViT-L/14-Reg was used as the backbone.
@@ 5.3 Cross-dataset detection
Table 8: Comparison of performance between various ConvNet and transformer architectures on the unseen test set, comprising images generated or manipulated by diffusion-based methods. In Approach 1, “blocks” denotes the number of final blocks utilized for feature extraction, while in Approach 2, “blocks” signifies the number of fine-tuned blocks.
@@
In this experiment, we assessed the generalizability of the detectors in detecting unseen deepfakes. The scenario presented a robust challenge, as there were no diffusion images in the training set. The classification thresholds were recalibrated using the unseen validation set. The results are presented in Table 8. Notably, there were drops in the performance of all models, with the best one going from 11.32% to 27.61% in terms of EER. Overall, Approach 2 consistently outperformed Approach 1. Within Approach 2, EfficientNetV2 exhibited better generalizability compared to other supervised pre-trained transformers. DINOv2 maintained its position as the top performer, reaffirming the superiority of using SSL with ViTs.
@@ 5.4 Visualization and explainability
With Approach 2, we can naturally visualize the focus areas of the DINO-based model using attention weights. To simplify the process, we computed the average of the attention maps from all attention heads directed toward the CLS token. The results are depicted in Fig. 3. To underscore the efficacy of fine-tuning, we compared the outcomes with those of the corresponding frozen original model. The partially fine-tuned model primarily directed its attention to the forehead, eyes, nose, and mouth to assess the authenticity of the input image. This behavior closely mirrors human intuition in deepfake detection, as deepfake artifacts frequently manifest in these regions. Notably, the original version of DINO did not possess this ability. Even when presented with unseen deepfakes, the fine-tuned model consistently prioritized these areas. This explains the model’s failure to detect deepfakes generated by Repaint-LDM, where the modification occurs in the hair region. In summary, such visualizations play a crucial role in deepfake detection, enhancing the interpretability of the results. The partially fine-tuned DINO model excelled in this regard.
@@ 6 Conclusion and future work
In this study, we explored two strategies for utilizing SSL pre-trained ViTs, specifically DINOs, as feature extractors for deepfake detection. The first approach involved utilizing frozen ViT backbones to extract multi-level features, while the second approach entailed partial fine-tuning on the final $k$ blocks. Through extensive experimentation, we found that the fine-tuning approach demonstrated superior performance and interpretability, particularly through attention mechanisms to visualize the focused areas. Our findings provide valuable insights for the digital forensic community regarding the utilization of SSL pre-trained ViTs as feature extractors, a relatively underexplored area in the literature of deepfake detection.
@@ Acknowledgements
This work was partially supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, and by JST CREST Grants JPMJCR18A6 and JPMJCR20D3, and by JST AIP Acceleration Grant JPMJCR24U3 Japan.

### v2

@@ 5.2 Approach 2: Fine-tuning final transformer blocks
We fine-tuned the last blocks (and the tokens in the case of transformers) of the selected backbones and compared their performances. Details are shown in Tab. 6. Compared to Approach 1, all backbones gained better results, with EVA-CLIP being the closest competitor with the SSL backbones. Nevertheless, DINOv2 and MAE remained the top performers. To narrow its gaps with DINOv2 and MAE, EVA-CLIP would need to be pre-trained with a vast dataset featuring rich annotations—a costly endeavor compared to DINOv2 and MAE, which was pre-trained on a substantially smaller dataset without any annotations. Given the same architecture (DeiT III versus DINOv2 and MAE), the performance gaps are significant, thanks to the training receipts. Overall, these results again underscore the significant advantage of using SSL for pre-training ViTs.
Next, we conducted an ablation study to determine the optimal number of $k$ final blocks required for fine-tuning. It is important to note that different ViT backbones have varying numbers of blocks; for example, DINOv2 - ViT-L/14-Reg has 24 blocks. The results are visualized in Fig. 2. If $k$ is small, the model may not adapt adequately to the new task, resulting in underfitting. Conversely, a large $k$ can lead to overfitting, especially with a small fine-tuning dataset. For DINOv2, the optimal $k$ is about half of the total blocks (11), while for MAE, it is three-fifths (15). The EERs decreased from 11.32% to 5.63% (-5.69%) for DINOv2 and from 10.34% to 6.88% (-3.46%) for MAE. When fine-tuning only the last block, MAE had about 1% better performance compared to DINOv2. However, with the optimal $k$ blocks, DINOv2 surpassed MAE, improving by about 2% compared to MAE’s 1%.
Figure 2: Ablation study on the relationship between the number of fine-tuned transformer blocks ($k$) and the EER in Approach 2.
Regarding the breakdown of results for both approaches, the real parts of FF++, Google DFD, DF-TIMIT, YouTube-DF, and DFDC are the most challenging subsets for detection. This difficulty may be due to the low quality of the deepfake media and diversity of deepfake methods in these subsets. Low quality can destroy artifacts, making detection harder. Fine-tuning with optimal $k$ blocks improved performance for Approach 2, but the real part of Google DFD still remained the most challenging.
@@ 5.3 Cross-dataset detection
Table 7: Comparison of performance between various ConvNet and transformer architectures on the unseen test set, comprising images generated or manipulated by diffusion-based methods. $k$ denotes the number of final blocks utilized for feature extraction in Approach 1 while signifies the number of fine-tuned blocks in Approach 2.
@@
DINOv2 ViT-L/14-Reg 11 0.7418 70.22 53.22 73.22 93.00 74.56 72.84 70.22 73.50 27.61 28.14
In this experiment, we assessed the generalizability of the detectors in detecting unseen deepfakes. The scenario presented a robust challenge, as there were no diffusion images in the training set. The classification thresholds were recalibrated using the unseen validation set. The results are presented in Tab. 7. Notably, there were drops in the performance of all models, with the best one going from 11.32% to 27.61% in terms of EER. Overall, Approach 2 consistently outperformed Approach 1. Within Approach 2, EfficientNetV2 exhibited better generalizability compared to other supervised pre-trained transformers.
DINOv2 is the absolute winner and clearly outperformed MAE. This can be explained by the SSL pre-training phase; DINOv2 employs strong data augmentations, while MAE uses little or none, making MAE less robust against unseen distributions. These results indicate that having the right SSL training strategy greatly enhances deepfake detection performance and is crucial for improving the generalizability of the backbone.
@@ 5.4 Visualization and explainability
With Approach 2, we can naturally visualize the focus areas of the ViT-based models using attention weights. To simplify the process, we computed the average of the attention maps from all attention heads directed toward the CLS token. We chose DINOv2 - ViT-L/14-Reg and randomly selected images per category for visualization to avoid cherry-picking. The results are depicted in Fig. 3. To highlight the efficacy of fine-tuning, we compared the outcomes with those of the corresponding frozen original model. The partially fine-tuned model primarily directed its attention to the forehead, eyes, nose, and mouth to assess the authenticity of the input image. This behavior closely mirrors human intuition in deepfake detection, as deepfake artifacts frequently manifest in these regions. Notably, the original version of DINO did not possess this ability. Even when presented with unseen deepfakes, the fine-tuned model consistently prioritized these areas. This explains the model’s failure to detect deepfakes generated by Repaint-LDM, where the modification occurs in the hair region. In summary, such visualizations play a crucial role in deepfake detection, enhancing the interpretability of the results. The partially fine-tuned DINO model excelled in this regard.
@@ 6 Conclusion and future work
In this study, we explored two strategies for utilizing SSL pre-trained ViTs–specifically DINOs and MAE–as feature extractors for deepfake detection. The first approach involved utilizing frozen ViT backbones to extract multi-level features, while the second approach entailed partial fine-tuning on the final $k$ blocks. Through extensive experimentation, we found that with a suitable SSL pre-training strategy, the fine-tuning approach demonstrated superior performance and interpretability, particularly through attention mechanisms to visualize the focused areas. Our findings provide valuable insights for the digital forensic community regarding the utilization of SSL pre-trained ViTs as feature extractors, a relatively underexplored area in the literature of deepfake detection.
@@ Acknowledgements
This work was partially supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.